Find the equation of the circle which has a diameter whose endpoints are (-6,3) and (2,5)
Answer by scott8148(3752)

the center of the circle is the midpoint of the diameter

(-6+2)/2 = -2

(3+5)/2 = 4 ___ so the center is (-2,4)

the radius is half the length of the diameter

sqrt[(2+6)^2 + (5-3)^2] = sqrt(68) = 2 sqrt(17)
so radius is sqrt(17)

using the general equation for a circle ___ (x+2)^2 + (y-4)^2 = 17